OpenStudy (anonymous):

Algebra help! Factor x^5+32

jhonyy9 (jhonyy9):

32 =2^5 ok ?

jhonyy9 (jhonyy9):

hope this will help you

OpenStudy (anonymous):

I understand that but I still dont know how to factor it

jhonyy9 (jhonyy9):

there are x^5 +2^5 so how can you factoring these ?

jhonyy9 (jhonyy9):

factorizing sorry

jhonyy9 (jhonyy9):

do you have any idea ?

jhonyy9 (jhonyy9):

x^5 =x^2 *x^3 yes ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would you factor is like a sum of squares and a sum of cubes?

jhonyy9 (jhonyy9):

yes right think about !!!

OpenStudy (anonymous):

(x^2+4)(x+2)(x^2-2x+4) is this right?

jhonyy9 (jhonyy9):

so i think you can check it too the same like me ,yes ?

OpenStudy (anonymous):

how?

jhonyy9 (jhonyy9):

(x^3 +2x^2 +4x +8)(x^2 -2x +4)= =x^5 -2x^4 +4x^3 +2x^4 -4x^3 +8x^2 +4x^3 -8x^2 +16x +8x^2 -16x +32= =x^5 +4x^3 +8x^2 +32 so than result that not is right

jhonyy9 (jhonyy9):

you have missed some factors there

OpenStudy (raden):

maybe we can use factor's theorema, to factorizing it because x=-2 is one of the root of x^5+32, so (x+2) is one of factor of x^5+32. just do synthetic division it will work

OpenStudy (anonymous):

can someone show me how? I dont know what factors theorema is or what synthetic division is.

OpenStudy (raden):